OpenStudy (anonymous):

Given A = {1, 2, 3, 4}, B = {4, 5, 6,}, and C = {2, 6, 7}. Evaluate each set a) A ∩ B b) A U C c) B U C d) (A U B) ∩ C e) A U (B U C) f) (A ∩ B) ∩ C g) (A ∩ B) U C

OpenStudy (anonymous):

So starting at the top: A intersect B is everything in both A and B.

OpenStudy (anonymous):

so it would be 4 thats in both

OpenStudy (anonymous):

Correct, {4} is the answer to question a)

OpenStudy (anonymous):

ok now the U i am a little confused on that one.

OpenStudy (anonymous):

is that everything that NOT in A and C?

OpenStudy (anonymous):

A union B is all the items that are in A or in B (or both)

OpenStudy (anonymous):

ok.. that makes sense..

OpenStudy (anonymous):

Oh, sorry the question is A union C

OpenStudy (anonymous):

So all the items in A, or C, or both.

OpenStudy (anonymous):

so the answer would be 4 cuz they both have it

OpenStudy (anonymous):

No. \(A\cup C\) = {1,2,3,4,6,7}

OpenStudy (anonymous):

so Union means ever element in both ones

OpenStudy (anonymous):

Because 1 is in A, 2 is in both, 3 is in A, 4 is in A, 6 is in C, 7 is in C.

OpenStudy (anonymous):

No. Union means all the elements in EITHER. Intersection means all the elements in BOTH.

OpenStudy (anonymous):

the upside down U is intersection and the U is union right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Try c)

OpenStudy (anonymous):

that would be {6} cuz they B and C have it only same

OpenStudy (anonymous):

B union C

OpenStudy (anonymous):

Not the ones in both, the ones in either

OpenStudy (anonymous):

wait B U C={2,4,5,6,7}

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Next one.

OpenStudy (anonymous):

First do A union B then take the intersection of that and C.

OpenStudy (anonymous):

A={1,2,3,4} B={4,5,6} (A U B={1,2,3,4,5,6} C={2,6,7} (A U B) ∩ C={2,6}

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

YES!!!!!! :) :) :) i think im finally getting this .. woo hooo.. lol

OpenStudy (anonymous):

I think you are also. Go for the next one.

OpenStudy (anonymous):

so heres a question. do ( ) first right? then do the rest of problem?

OpenStudy (anonymous):

Yep stuff in parens is evaluated first.

OpenStudy (anonymous):

ok working on problem.. give me minute

OpenStudy (anonymous):

) A U (B U C) B={4,5,6} C={2,6,7} (B U C)={2,4,5,6,7} A={1,2,3,4} answer A U (B U C)= {1,2,3,4,5,6,7}

OpenStudy (anonymous):

right? :)

OpenStudy (anonymous):

polpak?

OpenStudy (anonymous):

Sorry, one sec

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Yes, that is right

OpenStudy (anonymous):

YES!~ i think im finally getting this.. about damn time.. lol only taking college math for the 2nd time lol

OpenStudy (anonymous):

I've been there ;)

OpenStudy (anonymous):

it sucks.. and its a required class.. :(

OpenStudy (anonymous):

Oh, no. Math is great. Just keep at it. You're doing well

OpenStudy (anonymous):

e) A U (B U C) B={4,5,6} C={2,6,7} (B U C)={2,4,5,6,7} A={1,2,3,4} answer A U (B U C)={1,2,3,4,5,6,7}

OpenStudy (anonymous):

Isn't that the one you already did?

OpenStudy (anonymous):

no this is a different one.. lol

OpenStudy (anonymous):

lol yea it was the same.. my bad.. lol

OpenStudy (anonymous):

but good job doing it right twice in a row. means you definitely get it ;)

OpenStudy (anonymous):

lol ok so the upside down means everything that is the same right? just wanted to make sure i got this correct .. :)

OpenStudy (anonymous):

the intersection operator \(\cap\) means everything that is in BOTH. If it is just in one, or the other it doesn't count. It has to be in both.

OpenStudy (anonymous):

f) (A ∩ B) ∩ C A={1,2,3,4} B={4,5,6} (A ∩ B)={4} C={2,6,7} answer (A ∩ B) ∩ C={4}

OpenStudy (anonymous):

wait this is a trick one.. cuz there is nothing that is in both

OpenStudy (anonymous):

There is nothing that is common to A, B and C. So the intersection will be the empty set, {}

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Good catch

OpenStudy (anonymous):

g) (A ∩ B) U C A={1,2,3,4} B={4,5,6} (A ∩ B)={4} C={2,6,7} answer (A ∩ B) U C={2,4,6,7}

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

sweet! i finally got this type of problems down.. :) whoo hooo .. lmao Thank you for taking the time to teach me this stuff.. :)

OpenStudy (anonymous):

Of course! Happy to help =)

OpenStudy (anonymous):

Using correct mathematical notation and symbols, express the following in set-builder notation: Z = {8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18}

OpenStudy (anonymous):

im not sure how to do this ?

OpenStudy (anonymous):

It's trying to get you to write it in a more compact way.

OpenStudy (anonymous):

ok... i dont know how else to write it..lol

OpenStudy (anonymous):

Ok, there isn't a general way to explain set builder so I'll just show you and explain how to do this one and you'll hopefully get the idea. Z = \(\{ x | (x \in \mathbb{N} )\wedge (8 \le x \le 18)\} \)

OpenStudy (anonymous):

Do you know all those symbols?

OpenStudy (anonymous):

hmmmm no... :(

OpenStudy (anonymous):

Ok I'll list them: | means 'such that' \(\in\) means 'is an element of' \(\mathbb{N}\) is the set of natural numbers (1,2,3,4,... etc) \(\wedge\) means 'and' \(\le \) means 'less than or equal to' Using that translation, convert what I wrote into an english sentence. Z is the set of ...

OpenStudy (anonymous):

Just throw something out and I'll correct as needed.

OpenStudy (anonymous):

i am confused.. sorry

OpenStudy (anonymous):

what is x again? u dont have it listed

OpenStudy (anonymous):

x is just a variable called x

OpenStudy (anonymous):

i am very confused sorry.. :( on the whole thing..

OpenStudy (anonymous):

Ok I'll translate, you compare what I say to the symbol version as well as the original version and see if it makes sense: Original: Z = {8,9,10,11,12,13,14,15,16,17,18} Set builder: Z=\(\{ x | (x \in \mathbb{N} )\wedge (8 \le x \le 18)\}\) English: Z is the set of numbers x such that x is in the natural numbers and x is between 8 and 18 inclusive.

OpenStudy (anonymous):

so x would be 8,9,10,11,12,... ect

OpenStudy (anonymous):

such as the natural numbers

OpenStudy (anonymous):

ok i am starting to kinda understand this.. "kinda" lol

OpenStudy (anonymous):

Yes. It would have to be in order to match the original set.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

But we don't want all the natural numbers because then we'd have 1, 2, and 3, along with 3076, etc.

OpenStudy (anonymous):

so we say that x has to be in the natural numbers and has to be between 8 and 18.

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

But we have to say x is in the natural numbers. If we just said x was between 8 and 18 it would include numbers like 17.1254

OpenStudy (anonymous):

or 9.15

OpenStudy (anonymous):

we only want the 'whole' numbers

OpenStudy (anonymous):

ok now i understand it

OpenStudy (anonymous):

ty sweetie!

OpenStudy (anonymous):

Certainly

OpenStudy (anonymous):

:)